OpenStudy (anonymous):

Write the equation of the line which passes through (6 ,-6) and(8,-5) in slope intercept form

OpenStudy (tkhunny):

Have you considered the 2-point form?

OpenStudy (whpalmer4):

Find the slope via \[m = \frac{y_2-y_1}{x_2-x_1}\]Use point-slope form to write the equation of a line with slope \(m\) passing through a known point \((x_0,y_0)\): \[y-y_0 =m(x-x_0)\]and rearrange into slope-intercept form by adding \(y_0\) to both sides and simplifying.